OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. sin 52° cos 13° - cos 52° sin 13°? sin 39° sin 65° cos 65° cos 39°

OpenStudy (whpalmer4):

\[\sin(u-v) = \sin u\cos v -\cos u \sin v\]

OpenStudy (anonymous):

thank you

OpenStudy (whpalmer4):

what's your answer?